OpenStudy (anonymous):

How to solve for x in x+x+2+2x-4+x+2+2x=56

OpenStudy (anonymous):

x=8 right?

OpenStudy (anonymous):

x+x+2+2x-4+x+2+2x=56 2x+2+2x-4+x+2+2x= 2x+2x+x+2x=56 7x=56 x=8

OpenStudy (anonymous):

anything else ....feell free to ask. lol

OpenStudy (anonymous):

nope im good

OpenStudy (anonymous):

np...thanks